OpenStudy (anonymous):

Add 3x x ____ + _____ x^2 - 81 x-9

jimthompson5910 (jim_thompson5910):

\[\large \frac{3x}{x^2-81}+\frac{x}{x-9}\] \[\large \frac{3x}{(x-9)(x+9)}+\frac{x}{x-9}\] \[\large \frac{3x}{(x-9)(x+9)}+\frac{x(x+9)}{(x-9)(x+9)}\] \[\large \frac{3x}{(x-9)(x+9)}+\frac{x^2+9x}{(x-9)(x+9)}\] \[\large \frac{3x+x^2+9x}{(x-9)(x+9)}\] \[\large \frac{x^2+12x}{(x-9)(x+9)}\] \[\large \frac{x^2+12x}{x^2-81}\] So \[\large \frac{3x}{x^2-81}+\frac{x}{x-9}=\frac{x^2+12x}{x^2-81}\]